# A NICER view of the â edge-on pulsar PSR J06143329

### v1

@@ 6.4 Different modes and surface geometries
As the inferred distributions appear to have different modes whose distributions do not overlap at in some region of the parameter space, using the multi-modal version of MultiNest increases the acceptance rate of samples, which can help to reduce computation time while retaining sampling accuracy (Shaw et al., 2007; Feroz et al., 2009). The inferred radii differ for the two possible headline geometries (Figure 4), which individually have different radius values and lower uncertainties (see posterior distributions in Appendix B). Hence, breaking this multi-modality and converging on one single mode would reduce the total uncertainty in the radius measurement by a few hundred meters. Physically motivated hot region priors can help favoring one geometry over the other by ruling out geometric configurations (Kalapotharakos et al., 2021; Das et al., 2024). Such priors can also help differentiate between ST+PDTâ and PDT-Uâ models.
Future improvements in the available data with more NICER exposure, more accurate priors on distance, inclination and mass from radio pulsar timing (e.g., with the Square Kilometer Array, Basu et al., 2025) can also help break this multi-modality. Higher quality data obtained with next-generation X-ray observatories such as eXTP (, Li et al., 2025) and NewAthena (, Cruise et al., 2025) will also be of great help for this endeavor. These observatories will also decrease uncertainties of individual modes thanks to their large effective area, time resolution and constrained background, the latter being one of the main limitation of our study.
@@ 7 Conclusion
Various tests were carried out to assess the validity of different inferences. It was found that 1) the effect of LR likelihood on sampling for the exploratory runs was limited, 2) the inferred parameter posterior distributions were not affected by noise realization and 3) model comparison disfavored ST-Uâ and did not significantly favor PDT-Uâ over ST+PDT. As inferred radii and geometries were consistent, we chose the simpler ST+PDTâ model, which explains well the data, to run the Headline result.
@@
Further EOS analyses of the impact of this radius measurement result in a slight softening of the EOS compared to our previous work (Rutherford et al., 2024), as well as shifting the allowed mass-radius region toward lower radii by âm for both high-density EOS extensions. For the maximum mass, this new radius constraint has little impact.
This radius measurement could be narrowed down in the future with better geometry priors to break degeneracies, with improved X-ray data from XMM, NewAthena or eXTP to better constrain the background or with a tighter prior on the mass from radio timing.
Acknowledgments: The authors acknowledge fruitful discussion with Vivek Krishnan and Michael Kramer about the pulsar mass measurements from radio timing with MeerKAT. This study has been partially supported through the grant EUR TESS N°ANR-18-EURE-0018 in the framework of the Programme des Investissements dâAvenir. This work was performed using HPC resources from CALMIP (Grant 2016-19056). Part of this work has been supported by the CEFIPRA grant IFC/F5904-B/2018 and ANR-20-CE31-0010 (MORPHER). All researchers from IRAP acknowledge the support of the CNES. Portions of this work performed at NRL were supported by NASA. D.C., B.D. and A.L.W. acknowledge support from ERC Consolidator grant No. 865768 AEONS (PI: Watts). M.H., Y.K. and A.L.W. acknowledge support from the NWO grant ENW-XL OCENW.XL21.XL21.038 Probing the phase diagram of Quantum Chromodynamics (PI: Watts). The work of M.M., A.S., and I.S. was supported by the European Research Council (ERC) under the European Unionâs Horizon 2020 research and innovation program (Grant Agreement No. 101020842). N.R. acknowledges support from NASA grant No.80NSSC22K0092. S.B. acknowledges funding from NASA grants 80NSSC20K0275 and 80NSSC22K0728. The work presented is based in part on observations with XMM-Newton, an ESA Science Mission with instruments and contributions directly funded by ESA Member states and NASA. This research has made use of data products and software provided by the High Energy Astrophysics Science Archive Research Center (HEASARC), which is a service of the Astrophysics Science Division at NASA/GSFC and the High Energy Astrophysics Division of the Smithsonian Astrophysical Observatory. This work has relied on NASAâs Astrophysics Data System (ADS) bibliographic services and the ArXiv.
@@
Softwares: ArviZ (Kumar et al., 2019), Cython (Behnel et al., 2011), fgivenx (Handley, 2018), GetDist (Lewis, 2019, https://github.com/cmbant/getdist), GNU Scientific Library (GSL; Gough, 2009), HEASoft (Nasa High Energy Astrophysics Science Archive Research Center (2014), HEASARC), IPython (Perez & Granger, 2007), Jupyter (Kluyver et al., 2016), Matplotlib (Hunter, 2007; Droettboom et al., 2018), MPI (Forum, 1994), MPI for Python (DalcĂn et al., 2008), MultiNest (Feroz et al., 2009), NEOST v2.1.0 (https://github.com/xpsi-group/neost; Raaijmakers et al. 2025), nestcheck (Higson, 2018; Higson et al., 2018, 2019), NumPy (van der Walt et al., 2011), OpenMP (Dagum & Menon, 1998), PyMultiNest (Buchner et al., 2014), Python/C language (Oliphant, 2007), SAS (Gabriel et al., 2004) SciPy (Virtanen et al., 2020), X-PSI v2.2.7 (https://github.com/xpsi-group/xpsi; Riley et al. 2023)
@@ Appendix A Test of X-PSI resolution
The different settings whose effects on likelihood accuracy were studied are given in Table 3. sqrt_num_cells is the square root of the approximate number of cells used to represent a spot and num_rays is the number of emission angle grid points for which light bending angles are precomputed. When the hot region has two temperature components (ceding and superseding), min_sqrt_num_cells and max_sqrt_num_cells are respectively the minimum and maximum values of sqrt_num_cells for the individual components. num_leaves and num_energies are, respectively, the sizes of the regularly spaced arrays of phase and energy over which the model is computed. Folding with the instrument response and integrating over the data phase bins, we get a modeled observation for a set of parameters, which is compared to the data to compute the likelihood. Increasing any of these values improves the accuracy of the likelihood computation.
@@ Appendix B Headline result
The Headline ST+PDTâ run summary is given in Table 4. The posterior distribution of a subset of parameters from the headline ST+PDTârun, with individual modes detailed, is given in Figure 12. The figure of the full posterior distributions, including all parameters, can be found in the online version of this article.

### v2

@@ 6.4 Different modes and surface geometries
As the inferred distributions appear to have different modes whose distributions do not overlap at in some region of the parameter space, using the multi-modal version of MultiNest increases the acceptance rate of samples, which can help to reduce computation time while retaining sampling accuracy (shaw_efficient_2007; MultiNest_2009). The inferred radii differ for the two possible headline geometries (Figure 4), which individually have different radius values and lower uncertainties (see posterior distributions in Appendix B). Hence, breaking this multi-modality and converging on one single mode would reduce the total uncertainty in the radius measurement by a few hundred meters.
Physically motivated hot region priors can help favoring one geometry over the other by ruling out geometric configurations (Kalapotharakos_geometry_2009; Das_PPM_GRMHD_2025). Such priors can also help differentiate between ST+PDTâ and PDT-Uâ models.
Future improvements in the available data with more NICER exposure, more accurate priors on distance, inclination and mass from radio pulsar timing (e.g., with the Square Kilometer Array, ska_2025_densematter) can also help break this multi-modality. As the XMM exposure is only ks long888To put this in context with previous analyses, the MOS1/2 exposures were ks for PSRÂ J07406620 (salmi_radius_2022), ks long for PSRÂ J00300451 (vinciguerra_updated_2024), and ks long for PSRÂ J04374715 (choudhury_nicer_2024)., more XMM data would be valuable to further constrain the background, one of the main limitations of our study. This will in turn help to tighten the uncertainties on the inferred radius. Higher quality data obtained with next-generation X-ray observatories such as eXTP (, Li_eXTP_2025) and NewAthena (, cruise_newathena_2024) will also be of great help for this endeavor. These observatories will also decrease uncertainties of individual modes thanks to their large effective area, time resolution and constrained background.
@@ 7 Conclusion
Various tests were carried out to assess the validity of different inferences. It was found that (1) the effect of LR likelihood on sampling for the exploratory runs was limited, (2) the inferred parameter posterior distributions were not affected by noise realization, and (3) model comparison disfavored ST-Uâ and did not significantly favor PDT-Uâ over ST+PDT. As inferred radii and geometries were consistent, we chose the simpler ST+PDTâ model, which explains well the data, to run the Headline result.
@@
Further EOS analyses of the impact of this radius measurement result in a slight softening of the EOS compared to our previous work (rutherford_constraining_2024), as well as shifting the allowed mass-radius region toward lower radii by âm for both high-density EOS extensions, making them compatible with previous analyses within less than one standard deviation. For the maximum mass, this new radius constraint has little impact.
This radius measurement could be narrowed down in the future with better geometry priors to break degeneracies, with improved X-ray data from XMM, NewAthena, or eXTP to better constrain the background or with a tighter prior on the mass from radio timing.
Acknowledgments: The authors acknowledge fruitful discussion with Vivek Krishnan and Michael Kramer about the pulsar mass measurements from radio timing with MeerKAT. The authors thank the referee for their suggestions, which contributed to improve this article. This study has been partially supported through the grant EUR TESS N°ANR-18-EURE-0018 in the framework of the Programme des Investissements dâAvenir. This work was performed using HPC resources from CALMIP (Grant 2016-19056). Part of this work has been supported by the CEFIPRA grant IFC/F5904-B/2018 and ANR-20-CE31-0010 (MORPHER). All researchers from IRAP acknowledge the support of the CNES. Portions of this work performed at NRL were supported by NASA. D.C., B.D. and A.L.W. acknowledge support from ERC Consolidator grant No. 865768 AEONS (PI: Watts). M.H., Y.K. and A.L.W. acknowledge support from the NWO grant ENW-XL OCENW.XL21.XL21.038 Probing the phase diagram of Quantum Chromodynamics (PI: Watts). The work of M.M., A.S., and I.S. was supported by the European Research Council (ERC) under the European Unionâs Horizon 2020 research and innovation program (Grant Agreement No. 101020842). N.R. acknowledges support from NASA grant No.80NSSC22K0092. S.B. acknowledges funding from NASA grants 80NSSC20K0275 and 80NSSC22K0728. The work presented is based in part on observations with XMM-Newton, an ESA Science Mission with instruments and contributions directly funded by ESA Member states and NASA. This research has made use of data products and software provided by the High Energy Astrophysics Science Archive Research Center (HEASARC), which is a service of the Astrophysics Science Division at NASA/GSFC and the High Energy Astrophysics Division of the Smithsonian Astrophysical Observatory. This work has relied on NASAâs Astrophysics Data System (ADS) bibliographic services and the ArXiv.
@@
Softwares: ArviZ (kumar_arviz_2019), Cython (cython2011), fgivenx (fgivenx), GetDist (Lewis19, https://github.com/cmbant/getdist), GNU Scientific Library (GSL; Gough:2009), HEASoft (2014ascl.soft08004N), IPython (IPython2007), Jupyter (Kluyver:2016aa), Matplotlib (Hunter:2007; matplotlibv2), MPI (MPI), MPI for Python (mpi4py), MultiNest (MultiNest_2009), NEOST v2.1.0 (https://github.com/xpsi-group/neost; Raaijmakers24), nestcheck (higson2018nestcheck; higson2018sampling; higson2019diagnostic), NumPy (Numpy2011), OpenMP (openmp), PyMultiNest (PyMultiNest), Python/C language (python2007), SAS (2004ASPC..314..759G) SciPy (Scipy), X-PSI v2.2.7 (https://github.com/xpsi-group/xpsi; xpsi)
@@ Appendix A Test of X-PSI resolution
The different settings whose effects on likelihood accuracy were studied are given in Table 3. sqrt_num_cells is the square root of the approximate number of cells used to represent a spot and num_rays is the number of emission angle grid points for which light bending angles are precomputed. When the hot region has two temperature components (ceding and superseding), min_sqrt_num_cells and max_sqrt_num_cells are, respectively, the minimum and maximum values of sqrt_num_cells for the individual components. num_leaves and num_energies are, respectively, the sizes of the regularly spaced arrays of phase and energy over which the model is computed. Folding with the instrument response and integrating over the data phase bins, we get a modeled observation for a set of parameters, which is compared to the data to compute the likelihood. Increasing any of these values improves the accuracy of the likelihood computation.
@@ Appendix B Headline result
The Headline ST+PDTâ run summary is given in Table 4. The posterior distribution of a subset of parameters from the headline ST+PDTârun, with individual modes detailed, is given in Figure 12. The figure of the full posterior distributions, including all parameters, can be found in the Zenodo repository (zenodo).